OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal? Data set Mean Mean absolute deviation 1 10.7 2.4 2 13.7 2.5

OpenStudy (anonymous):

@jigglypuff314 @jhonyy9

OpenStudy (anonymous):

is it 1.6?

jhonyy9 (jhonyy9):

how you can calculi it the ratio ?

OpenStudy (anonymous):

well you first start off finding the mean and the MAD... afterwards u subtract it and the number subtracted is divided by the MAD... then do you add it?

jhonyy9 (jhonyy9):

why so complex ? what mean ratio of two numbers ? for example ratio of 10 and 5 equal how many ?

OpenStudy (anonymous):

it is 2

OpenStudy (anonymous):

but that is what the steps said

jhonyy9 (jhonyy9):

exactly - so make it same in case of your exercise too

OpenStudy (anonymous):

??? so do i do 10.7/2.4 and 13.7/ 2.5?

jhonyy9 (jhonyy9):

case 1- 10,7/2,4 = ? case 2 yes exactly

OpenStudy (anonymous):

so it is 4.46

jhonyy9 (jhonyy9):

every two ?

OpenStudy (anonymous):

no the 10.7/2.4

OpenStudy (anonymous):

13.7/2.5=5.48

jhonyy9 (jhonyy9):

ok, and for second ?

jhonyy9 (jhonyy9):

ok and now how you get the ratio result ?

jhonyy9 (jhonyy9):

assum these two and divide by two

OpenStudy (anonymous):

do we do 4.46/5.48? or 5.48/4.46?

jhonyy9 (jhonyy9):

not is the question the ratio of the ratio just ratio of means to MAD - yes ?

OpenStudy (anonymous):

... yeah

jhonyy9 (jhonyy9):

so from 1 step you have got 4,46 and for 2 step you have got 5,48 and from these what will be the resulted ratio ? assum these two and divide by 2 and will get the right ratio - ok. ?

OpenStudy (anonymous):

so if i divide 4.46 and 5.48, answer is 0.82?

OpenStudy (anonymous):

i mean 0.81

jhonyy9 (jhonyy9):

do you have any choices ?

OpenStudy (anonymous):

nope

jhonyy9 (jhonyy9):

for right answer ?

jhonyy9 (jhonyy9):

so than please read what i have wrote to you above

OpenStudy (anonymous):

oh...so 4.46+5.48=9.94/2=4.97?

jhonyy9 (jhonyy9):

yes i think this is the right way and answer

OpenStudy (anonymous):

ok thanks!

jhonyy9 (jhonyy9):

was my pleasure but do you understand it why so ?

OpenStudy (anonymous):

yes! thank you!

jhonyy9 (jhonyy9):

ok anytime bye bye

OpenStudy (anonymous):

the answer is actually 1.2